OpenStudy (anonymous):

What is an equation for the line with slope 2 over 3 and y-intercept 9? explain please

OpenStudy (anonymous):

You have a linear problem, you can write it in the form y =Ax + B Where A is the slope and B the y-intercept

OpenStudy (anonymous):

the equation is y=(2/3)x+9, in standard form it would be 2x-3y=-27